OpenStudy (anonymous):

Simplify the given expression: 4 divided by the quantity of 3 minus 2i

OpenStudy (anonymous):

Btw this operating complex numbers, really need help thanks !

OpenStudy (dumbcow):

get rid of imaginary numbers in denominator by multiplying by conjugate \[\frac{4}{(3-2i)} *\frac{(3 + 2i)}{(3+2i)}\]

OpenStudy (anonymous):

so now its 4/(3+2) ?

OpenStudy (anonymous):

@dumbcow

OpenStudy (solomonzelman):

you are getting the denominator correctly, but check your numerator.

OpenStudy (anonymous):

does it become negative? @SolomonZelman

OpenStudy (solomonzelman):

Does `what become negative ?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

the numerator becomes 16 right?

OpenStudy (solomonzelman):

\[4(3+2i) =?\]

OpenStudy (anonymous):

so 4(3) is 12 and 4(2i) is -8? so -4?

OpenStudy (anonymous):

The denominator should be 3^2 + 2^2 = 13

OpenStudy (anonymous):

I believe the final asnwer is 12-8i/13 is that correct>

OpenStudy (anonymous):

And the numerator = 4*(3+2i) = 12+8i Final result is (12+8i) / 13

OpenStudy (anonymous):

Ok thanks alot I actually am starting to understand :)